OpenStudy (anonymous):

Suppose that you roll a pair of fair dice. What is the probability that you roll a sum of 7 two times in a row?

OpenStudy (anonymous):

a sum of 7 can be obtained in six ways on each toss so,

OpenStudy (anonymous):

i get that the sum of 7 can be obtained in six ways. but why did you multiply the 1/6 by 1/6? how did you get that?

OpenStudy (nurali):

What is the probability that you roll a sum of 7 two times in a row Means there are total 6 die probability of 1 times die is 1/6 probibility of two times dice means (1/6)^2=1/36 sum of 7 two times means (1/6)^2=1/36

OpenStudy (anonymous):

oooh ok got it, thx so much!

OpenStudy (anonymous):

LOL